An individual who has automobile insurance from a certain company is randomly selected. Let Y be the number of moving violations for which the individual was cited during the last 3 years. The pmf of Y is the following. p(y) 0.60 0.25 0.10 0.05 (a) Compute E(Y). E(Y) (b) Suppose an individual with γ violations incurs a surcharge of $110. Calculate the expected amount of the surcharge.
Answer #1

a) E(Y) = 0*0.6 + 1*0.25 + 2*0.1 + 3*0.05 = 0.25 + 0.2 + 0.15 = 0.6

b) Surcharge per Y violations = 110*Y

E(Surcharge) = E(Y)*110

= 0.6*110

= 66

Hope this helps and Please don't forget to rate the answer :)